# Stoichiometry Representation Learning with Polymorphic Crystal Structures

### v1

@@ F.2 Transfer Learning
Supervised Learning
Rand init. 0.390 0.599 0.204 0.849 0.202 0.425 0.659 0.209 0.402
(0.012) (0.053) (0.014) (0.174) (0.027) (0.048) (0.098) (0.019) (0.082)
Roost 0.384 0.743 0.199 0.851 0.216 0.406 0.684 0.240 0.402
(0.008) (0.069) (0.023) (0.126) (0.037) (0.046) (0.180) (0.048) (0.054)
CrabNet 0.427 0.759 0.253 1.206 0.332 0.576 0.868 0.353 0.691
(0.012) (0.052) (0.023) (0.042) (0.046) (0.080) (0.138) (0.040) (0.057)
Transfer Learning
GraphCL 0.391 0.607 0.193 0.862 0.198 0.412 0.643 0.205 0.412
(0.011) (0.026) (0.018) (0.236) (0.031) (0.006) (0.098) (0.021) (0.077)
MP Band G. 0.382 0.604 0.193 0.829 0.210 0.405 0.632 0.197 0.402
(0.012) (0.036) (0.025) (0.187) (0.038) (0.006) (0.095) (0.028) (0.081)
MP Form. E. 0.391 0.582 0.197 0.822 0.195 0.410 0.641 0.209 0.428
(0.013) (0.015) (0.019) (0.167) (0.031) (0.041) (0.102) (0.043) (0.086)
3D Infomax 0.391 0.606 0.194 0.844 0.210 0.402 0.633 0.207 0.391
(0.006) (0.027) (0.019) (0.195) (0.032) (0.005) (0.133) (0.018) (0.077)
PolySRL 0.386 0.576 0.191 0.822 0.189 0.386 0.626 0.195 0.390
(0.021) (0.042) (0.024) (0.162) (0.037) (0.069) (0.161) (0.015) (0.077)
F.3 Physical Validity
Further Analysis. In this section, we delve deeper into the physical validity of predicted properties for thermoelectrical materials by observing scatter plots that compare the actual ground truth values of $Z\bar{T}$ with the values obtained by the model predictions. For clearer visualization, we select one baseline model from models that consider DFT-calculated properties (i.e., MP Band G.) and structures (i.e., 3D Infomax). In Figure 6, we notice that the predictions produced by PolySRL consistently yield accurate calculations of $Z\bar{T}$ without any outliers. This observation underscores the model’s ability to predict physically valid properties for thermoelectrical materials. Additionally, we observe that the model, specifically MP Band G., which lacks consideration of the structural information within stoichiometry, tends to produce outliers more frequently when contrasted with models that incorporate structural information. More specifically, three outliers made by MP Band G. in Figure 6 (a) are $\text{Co}_{9}\text{S}_{8}$, $\text{Cu}_{5}\text{Sn}_{2}\text{S}_{6.65}\text{Cl}_{0.35}$, and $\text{Cu}_{5.133}\text{Sn}_{1.866}\text{S}_{6.65}\text{Cl}_{0.35}$. In case of $\text{Co}_{9}\text{S}_{8}$, there exist only one possible structure in MP dataset, and there was no existing structure for $\text{Cu}_{5}\text{Sn}_{2}\text{S}_{6.65}\text{Cl}_{0.35}$, and $\text{Cu}_{5.133}\text{Sn}_{1.866}\text{S}_{6.65}\text{Cl}_{0.35}$. This suggests that MP Band G. encounters difficulty in acquiring accurate physical properties for materials where obtaining structural information is computationally challenging. On the other hand, in Figure 6 (b), two outliers made by MP Band G. are GeTe and SnTe, each of which has three possible structures in MP dataset. This indicates that MP Band G. suffers from obtaining valid physical properties from polymorphic structures. In conclusion, we argue that this finding underscores the significance of incorporating structural information for accurate predictions.
Figure 6: Scatter plot between true and predicted $Z\bar{T}$.
High Throughput Screening. As described in the main manuscript, the figure of merit $Z\bar{T}$ determines how effectively power can be generated and energy can be harvested across various real-world applications. To discover novel materials of high $Z\bar{T}$, we perform high-throughput screening based on the predicted $Z\bar{T}$ in Figure 7. In particular, for thermoelectrical materials at room temperature (300 K), we establish a threshold of $Z\bar{T}=0.8$, and for high-temperature scenarios (600 K), we use a threshold of $Z\bar{T}=1.1$. We observe that PolySRL outperforms all other baseline methods in ESTM 300K datasets, while performing competitively with 3D Infomax in ESTM 600K. This again demonstrates the importance of structural information in stoichiometry representation learning, which has been overlooked in previous works Goodall & Lee (2020); Wang et al. (2021).
Figure 7: High throughput screening results.
F.4 Model Analysis
@@
Ablation Studies. In this section, we conduct ablation studies on our model by removing the sampling process described in Equation 3, which is denoted as "w/o Sampling" in Table 5. To clarify, rather than utilizing the sampled representations $\hat{\textbf{z}}_{j}^{a}$ in Equation 4, we directly employ the mean vector of stoichiometry, denoted as $\textbf{z}_{\mu}^{a}$, for the soft contrastive loss. By doing so, the model transitions from learning a probabilistic representation of stoichiometry to learning a deterministic representation of stoichiometry. To compare with methods that don’t incorporate polymorphic structural information, such as 3D Infomax, we also present the performance of 3D Infomax in Table 5. We have the following observations: 1) Considering polymorphic structure is crucial in stoichiometry representation learning by comparing 3D Infomax and w/o Sampling. 2) Additionally, the sampling process typically leads to improved performance, underscoring the advantage of learning a probabilistic representation of stoichiometry. While w/o Sampling outperforms PolySRL in two datasets, the absence of the sampling process means the model can no longer estimate uncertainty in stoichiometry, thereby losing its practicality in real-world materials discovery. In summary, we argue that PolySRL learns a probabilistic stoichiometry representation, which not only enables accurate uncertainty estimation but also enhances model performance.
@@
Sensitivity Analysis on $\beta$. In this section, we verify the empirical effect of the hyperparameter $\beta$, which controls the weight of the KL divergence loss computed between the learned distributions and the standard normal distribution, in Equation 6. We have the following observations from Figure 8 (a): 1) As the hyperparameter $\beta$ increases, the average variance of the learned distributions (i.e., uncertainty) also increases, and the dimension of the variance vectors that collapse to zero (i.e., collapsed ratio) decreases. This indicates that the KL divergence loss effectively prevents the distributions from collapsing. 2) On the other hand, the performance of PolySRL deteriorates as $\beta$ increases, indicating that emphasizing the KL divergence loss too much causes PolySRL to struggle in learning high-quality stoichiometry representations. However, reducing $\beta$ does not always result in improved performance, as collapsed distribution may not effectively capture information from polymorphic structures. Hence, selecting an appropriate value of $\beta$ is vital for learning high-quality stoichiometry representations while maintaining a suitable level of uncertainty.
Sensitivity Analysis on Various Hyperparameters. In addition, we provide an analysis on various hyperparameters in PolySRL, i.e., initial values of $c,d$ and number of samples $J$ in Equation 5. We have the following observations: 1) While we made $c$ and $d$ learnable parameters to allow the model to adjust them adaptively to an optimal point, we’ve also found that setting the initial values for $c$ and $d$ is crucial in model training. This indicates that initial value plays a significant role in guiding the model correctly from the outset of the training process, ultimately contributing to good performance. 2) On the other hand, we observe PolySRL shows robustness in various numbers of samples, suggesting that it can be trained effectively without a large number of samples, which will demand an extensive amount of computational resources.
Figure 8: Additional sensitivity analysis results.
F.5 Additional Uncertainty Analysis
Uncertainty and Model Performance. In this section, we analyze how the model performance varies regarding the uncertainties of the stoichiometry. To achieve this, we initially categorize stoichiometry based on MAE into intervals such as 0.0 to 1.0, 1.0 to 2.0, $\cdots$, and 4.0 to 5.0. For example, Group 1 in Figure 9 (a) contains the group of MAE in the range 0.0 to 1.0. We then calculate the average uncertainties of the model for each group. As observed in Figure 9 (a), as the MAE values increase, the level of uncertainty also increases, demonstrating that the model effectively estimates uncertainties associated with MAE values.
@@ F.5 Additional Uncertainty Analysis
Additional Case Studies: Low Uncertainty with Multiple Structures. In addition to the case studies in Section 5.3, we further provide cases where the stoichiometry with multiple possible structures exhibits low uncertainty. In Figure 9 (b), we observe two stoichiometries with collapsed uncertainty, even though they possess four distinct possible structures. This phenomenon occurs because these structures share highly similar polymorphic arrangements, with only one unique structure in each stoichiometry. For instance, ZrC and $\text{NdF}_{2}$ predominantly adopt cubic and hexagonal structures, respectively, with only one distinct possible structure for each stoichiometry.
@@
Additional Case Studies: High Uncertainty with Multiple Structures. In this section, we present additional case studies that align with our expectations. Figure 9 (c) illustrates two stoichiometries with the highest uncertainty among those possessing three polymorphic structures. For example, NaI can exist in three distinct structures (i.e., cubic, orthorhombic, and tetragonal), and AlP also exhibits three different structures (i.e., cubic, hexagonal, and tetragonal). Given that varying atomic arrangements within materials lead to entirely distinct physical and chemical properties, it becomes crucial to convey the extent of structural diversity that stoichiometry can exhibit during the material discovery process. Therefore, these additional case studies highlight the practicality of PolySRL in real-world material discovery.
Figure 9: Additional uncertainty analysis.
Appendix G Notations
@@
Notations
Explanations
$n_{s}$
Number of atoms in crystal structure
$\mathbf{X}^{b}$
An elemental feature matrix of structural graph
$\mathbf{A}^{b}$
An adjacency matrix of structural graph
$\mathcal{G}^{b}=(\mathbf{X}^{b},\mathbf{A}^{b})$
A crystal structural graph
$\mathbf{z}^{b}$
A latent representation of a crystal structural graph
$f^{b}$
A GNN-based crystal structural encoder
$n_{e}$
Number of unique elements in a stoichiometry
=$\mathcal{E}=\{e_{1},\ldots,e_{n_{e}}\}$
A unique set of elements in a stoichiometry
=$\mathcal{R}=\{r_{1},\ldots,r_{n_{e}}\}$
A compositional ratio of each element in a stoichiometry
$\mathcal{G}^{a}=(\mathcal{E},\mathcal{R},\mathbf{A}^{a})$
A fully-connected stoichiometry graph
$\mathbf{X}^{a}$
A elemental feature matrix of stoichiometry graph
$\mathbf{A}^{a}$
An adjacency matrix of stoichiometry graph
$\tilde{\mathbf{z}}^{a}$
A sampled representation from latent distribution of stoichiometry
$f^{a}$
A GNN-based stoichiometry graph encoder
$f_{\mu}^{a}$
A mean module for stoichiometry graph
$f_{\sigma}^{a}$
A variance module for stoichiometry graph
$J$
Number of samples from latent distribution of stoichiometry (Equation 4)
$c$
Learnable parameters for scaling the Euclidean distance (Equation 4)
$d$
Learnable parameters for shifting the Euclidean distance (Equation 4)
$\mathcal{L}_{\text{con}}$
Soft contrastive loss (Equation 5)
$\mathcal{L}_{\text{KL}}$
KL divergence loss
$\beta$
Hyperparameter that controls the weight of KL divergence loss
$\mathcal{L}_{\text{total}}$
Total loss function (Equation 6)

### v2

@@ C.2 Statistical Significance Test
GraphCL $3.80e^{-4}$ $4.18e^{-3}$ $7.63e^{-3}$ $3.12e^{-3}$ $4.29e^{-4}$ $2.83e^{-3}$ $6.72e^{-4}$ $1.39e^{-2}$ $4.07e^{-2}$
MP Band G. $2.48e^{-1}$ $1.20e^{-4}$ $6.53e^{-4}$ $4.48e^{-4}$ $2.15e^{-4}$ $4.55e^{-4}$ $5.90e^{-4}$ $1.47e^{-3}$ $1.40e^{-3}$
MP Form. E. $5.14e^{-2}$ $9.23e^{-1}$ $1.89e^{-2}$ $3.93e^{-4}$ $1.75e^{-3}$ $5.03e^{-3}$ $1.47e^{-3}$ $3.23e^{-3}$ $4.30e^{-2}$
3D Infomax $1.21e^{-3}$ $1.40e^{-3}$ $3.42e^{-2}$ $1.44e^{-2}$ $2.72e^{-2}$ $8.39e^{-4}$ $6.06e^{-3}$ $3.89e^{-2}$ $5.64e^{-2}$
We have the following observations: 1) PCRL achieves statistically significant improvements over baseline methods that do not leverage structural information (such as GraphCL, MP Band G., MP Form. E.) at the 0.05 significance level except for Band G. and Form. E. datasets. 2) On the other hand, in the datasets for Band G. and Form. E., PCRL does not achieve statistical significance compared to models that benefit from the corresponding DFT-calculated values. For example, in Band G. dataset, we observe that the model trained with DFT-calculated band gap (i.e., MP Band G.) outperforms PCRL, which was also described in observation 3 in Section 3.1 in the main paper. However, as discussed in the main text, these models are specifically tailored for particular properties, which may limit their broader applicability across a range of tasks.
C.3 Fine-tuning on Out-of-Distribution Scenarios
In this section, we assess model performance in out-of-distribution scenarios by dividing the evaluation dataset according to composition information. Although we randomly split the dataset and conducted 5-fold cross-validation in the main manuscript, for this particular experiment, the training set excludes elements from the Actinide and Lanthanide series, whereas the test set includes them [74, 75]. The model parameters are fine-tuned with the training set and evaluated on the test set. In Table 9, we observe that PCRL surpasses baseline models in seven out of nine datasets, illustrating that the PCRL training method is also effective in out-of-distribution scenarios, which is crucial in real-world materials discovery.
Table 9: Fine-tuning performance on unseen types of materials (MAE).
Model DFT Poly. Band G. Form. E. Metallic ESTM 300K ESTM 600K
Prop. Str. E.C. T.C. Seebeck E.C. T.C. Seebeck
Rand init. 0.599 1.507 0.259 1.379 0.371 0.416 0.832 0.275 0.866
GraphCL 0.589 1.531 0.253 1.516 0.312 0.382 1.096 0.291 0.829
MP Band G. 0.563 1.239 0.262 1.757 0.421 0.556 0.836 0.302 0.975
MP Form. E. 0.570 1.318 0.259 1.580 0.397 0.454 0.993 0.247 0.780
3D Infomax 0.562 1.343 0.250 1.668 0.355 0.399 0.853 0.279 0.858
PCRL 0.546 1.186 0.244 1.320 0.412 0.403 0.794 0.236 0.729
C.4 Model Analysis
Sensitivity Analysis. In addition to model analysis in Section 3.1, we provide an analysis on various hyperparameters in PCRL, i.e., initial values of $c,d$ and number of samples $J$ in Equation 5. We have the following observations: 1) While we made $c$ and $d$ learnable parameters to allow the model to adjust them adaptively to an optimal point, we’ve also found that setting the initial values for $c$ and $d$ is crucial in model training. This indicates that initial value plays a significant role in guiding the model correctly from the outset of the training process, ultimately contributing to good performance. 2) On the other hand, we observe PCRL shows robustness in various numbers of samples, suggesting that it can be trained effectively without a large number of samples, which will demand an extensive amount of computational resources.
Figure 9: Additional sensitivity analysis results.
@@ C.4 Model Analysis
Ablation Studies. In this section, we conduct ablation studies on our model by removing the sampling process described in Equation 3, which is denoted as "w/o Sampling" in Table 10. To clarify, rather than utilizing the sampled representations $\hat{\textbf{z}}_{j}^{a}$ in Equation 4, we directly employ the mean vector of composition, denoted as $\textbf{z}_{\mu}^{a}$, for the soft contrastive loss. By doing so, the model transitions from learning a probabilistic representation of composition to learning a deterministic representation of composition. To compare with methods that don’t incorporate polymorphic structural information, such as 3D Infomax, we also present the performance of 3D Infomax in Table 10. We have the following observations: 1) Considering polymorphic structure is crucial in compositional representation learning by comparing 3D Infomax and w/o Sampling. 2) Additionally, the sampling process typically leads to improved performance, underscoring the advantage of learning a probabilistic representation of composition. While w/o Sampling outperforms PCRL in two datasets, the absence of the sampling process means the model can no longer estimate uncertainty in composition, thereby losing its practicality in real-world materials discovery. In summary, we argue that PCRL learns a probabilistic compositional representation, which not only enables accurate uncertainty estimation but also enhances model performance.
@@
C.5 Pre-training on OQMD Dataset
In Table 11, we present the comprehensive experimental results from Table 3.1 in the main manuscript. We observe that the model trained with the MP database generally performs better than the one trained with the OQMD database.
Table 11: Representation learning performance (MAE) comparison trained with different databases.
Database Band G. Form. E. Metallic ESTM 300 K ESTM 600 K
E.C. T.C. Seebeck E.C. T.C. Seebeck
MP (Ours) 0.407 0.592 0.194 0.912 0.197 0.388 0.665 0.189 0.412
(0.013) (0.039) (0.017) (0.121) (0.020) (0.059) (0.126) (0.017) (0.043)
OQMD 0.425 0.663 0.194 0.814 0.216 0.401 0.743 0.209 0.439
(0.013) (0.039) (0.017) (0.121) (0.020) (0.059) (0.126) (0.017) (0.043)
C.6 Many-to-many relationship between composition and structure
In Table 12, we present the comprehensive experimental results from Table 3.1 in the main manuscript. It is evident that the model which focuses on learning the one-to-many relationship between composition and structure typically surpasses the performance of the model that addresses the many-to-many relationship.
Table 12: Representation learning performance (MAE) comparison trained with different relationship.
Relationship Band G. Form. E. Metallic ESTM 300 K ESTM 600 K
E.C. T.C. Seebeck E.C. T.C. Seebeck
One-to-Many (Ours) 0.407 0.592 0.194 0.912 0.197 0.388 0.665 0.189 0.412
(0.013) (0.039) (0.017) (0.121) (0.020) (0.059) (0.126) (0.017) (0.043)
Many-to-Many 0.410 0.617 0.210 0.892 0.199 0.416 0.670 0.199 0.440
(0.012) (0.025) (0.023) (0.117) (0.027) (0.070) (0.110) (0.053) (0.066)
C.7 Case Studies
From now on, we show case studies that are in line with our uncertainty analysis in Section 3.2.
@@ C.7 Case Studies
Low Uncertainty with Multiple Structures. In Figure 10 (a), we observe two compositions (i.e., ZrC and $\text{NdF}_{3}$) with collapsed uncertainty, even though they possess four distinct possible structures. This phenomenon occurs because these structures share highly similar polymorphic arrangements, with only one unique structure in each composition. For instance, ZrC and $\text{NdF}_{3}$ predominantly adopt cubic and hexagonal structures, respectively, with only one distinct possible structure for each composition.
@@
High Uncertainty with Multiple Structures. In this section, we present additional case studies that align with our expectations. Figure 10 (b) illustrates two compositions (i.e., NaI and AlP) with the highest uncertainty among those possessing three polymorphic structures. For example, NaI can exist in three distinct structures (i.e., cubic, orthorhombic, and tetragonal), and AlP also exhibits three different structures (i.e., cubic, hexagonal, and tetragonal). Given that varying atomic arrangements within materials lead to entirely distinct physical and chemical properties, it becomes crucial to convey the extent of structural diversity that composition can exhibit during the material discovery process. Therefore, these additional case studies highlight the practicality of PCRL in real-world material discovery.
Figure 10: Additional qualitative uncertainty analysis.
Appendix D Related Works
D.1 Probabilistic Representation Learning
First appearing in 2014 with the introduction of probabilistic word embeddings [76], probabilistic representations got a surge of interest from ML researchers by offering numerous benefits in modeling uncertainty pertaining to a representation. Specifically, in the computer vision domain, [77] proposes to probabilistically represent face images to address feature ambiguity in real-world face recognition. Moreover, [21] introduces Hedged Instance Embeddings (HIB), which computes a match probability between point estimates but integrates it over the predicted distributions via Monte Carlo estimation. This idea has been successfully extended to cross-modal retrieval [22], video representation learning [78], and concept prediction [79]. On the other hand, the importance of measuring uncertainties goes beyond the field of computer vision, particularly in the context of applying machine learning in the scientific discovery process [80, 2, 1]. In our study, our objective is to develop a probabilistic representation of composition that incorporates uncertainties related to the different polymorphic forms associated with a single composition, thereby enhancing the reliability of the ML model for the material discovery process.
D.2 Crystal Structure Prediction
Crystal structure prediction (CSP) is the process of determining the stable three-dimensional structure of a compound from its chemical composition alone. Conventional CSP methods often combine density functional theory (DFT) with optimization algorithms. These algorithms carry out an iterative process searching for stable states that align with the local energy minima, while DFT is utilized to assess the energy at each step of the iteration [81, 82, 83]. A recent development in the field is DiffCSP [84], which employs a deep generative model within a diffusion framework to simultaneously optimize lattice matrices and atomic coordinates, offering a novel approach to the CSP problem.
It’s important to highlight the growing interest among researchers in generating crystal structures without predefined composition, a direction distinct from CSP. The groundbreaking CDVAE [85] integrates a variational autoencoder (VAE) architecture with a diffusion-based decoder to produce the types of atoms, their coordinates, and lattice parameters. Unlike CSP-focused methods, CDVAE’s primary goal is the generation of random crystal structures, providing a different avenue in the field of crystallography research.
Appendix E Notations
@@
Notations Explanations
$n_{s}$ Number of atoms in crystal structure
$\mathbf{X}^{b}$ An elemental feature matrix of structural graph
$\mathbf{A}^{b}$ An adjacency matrix of structural graph
$\mathcal{G}^{b}=(\mathbf{X}^{b},\mathbf{A}^{b})$ A crystal structural graph
$\mathbf{z}^{b}$ A latent representation of a crystal structural graph
$f^{b}$ A GNN-based crystal structural encoder
$n_{e}$ Number of unique elements in a composition
=$\mathcal{E}=\{e_{1},\ldots,e_{n_{e}}\}$ A unique set of elements in a composition
=$\mathcal{R}=\{r_{1},\ldots,r_{n_{e}}\}$ A compositional ratio of each element in a composition
$\mathcal{G}^{a}=(\mathcal{E},\mathcal{R},\mathbf{A}^{a})$ A fully-connected composition graph
$\mathbf{X}^{a}$ A elemental feature matrix of composition graph
$\mathbf{A}^{a}$ An adjacency matrix of composition graph
$\tilde{\mathbf{z}}^{a}$ A sampled representation from latent distribution of composition
$f^{a}$ A GNN-based composition graph encoder
$f_{\mu}^{a}$ A mean module for composition graph
$f_{\sigma}^{a}$ A variance module for composition graph
$\mathcal{P}^{\mathcal{G}^{a}}$ A set of polymorphic structural graphs related to composition $\mathcal{G}^{a}$
$J$ Number of samples from latent distribution of composition (Equation 4)
$c$ Learnable parameters for scaling the Euclidean distance (Equation 4)
$d$ Learnable parameters for shifting the Euclidean distance (Equation 4)
$\mathcal{L}_{\text{con}}$ Soft contrastive loss (Equation 5)
$\mathcal{L}_{\text{KL}}$ KL divergence loss
$\beta$ Hyperparameter that controls the weight of KL divergence loss
$\mathcal{L}_{\text{total}}$ Total loss function (Equation 6)